20160126 » History » Version 20
Boyd, Steven, 27 January 2016 08:14
1 | 1 | Boyd, Steven | h1. Commissioning and Ops Scheduling meeting : 26/1/2016 |
---|---|---|---|
2 | |||
3 | To be held in CR10 in the ATLAS Building @ 15:30 pm |
||
4 | |||
5 | Phone access : ReadyTalk number |
||
6 | US 866-740-1260 |
||
7 | UK 0800-496-0576 |
||
8 | Swiss 0800-70-53-51 |
||
9 | Italy 800182592 |
||
10 | Access Code: 2155402 |
||
11 | |||
12 | h2. Agenda: |
||
13 | |||
14 | * Review of tasks and activities identified at last meeting |
||
15 | ** Realignment work |
||
16 | 8 | Boyd, Steven | *** *Action Item (Colin to prod John C)* : need to get John's data with fiducials in it to get to Tony Middleton. _Expected yesterday_ |
17 | 1 | Boyd, Steven | ** Simulation of material in the interspace |
18 | *** *Action Item* : John to give these files to the tracker group to make certain tracker resolution doesn't change |
||
19 | *** *Action Item* : John to do a simulation with air in the interspace and He in the absorber volume |
||
20 | *** *Action Item* : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded |
||
21 | *** *Action Item* : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum |
||
22 | ** Cooling water discussion |
||
23 | 9 | Boyd, Steven | *** *Action Item* : Colin and John to talk and agree on the plan. _Discussion complete but further work needed_ |
24 | 1 | Boyd, Steven | ** ISIS Monitoring of SS magnets |
25 | *** *Action Item* : Josef/Colin to check with Alan about monitoring requirements for the SS magnets. |
||
26 | ** BLOCS and SOCS |
||
27 | 2 | Boyd, Steven | *** *Action Item* : -Steve to remind Adam that we need SOCs.- |
28 | 1 | Boyd, Steven | ** DS PS Installation |
29 | 8 | Boyd, Steven | *** *Action Item* : Colin is to get an estimate for power supply installation time. _2 sets of 2-3 days over a space of 2 weeks_ |
30 | 1 | Boyd, Steven | * Upcoming and ongoing activities : chaired by DC |
31 | ** "Operations calendar":https://www.google.com/calendar/embed?src=miceops%40gmail.com&ctz=Europe/London |
||
32 | ** C&M Network reconfiguration |
||
33 | 14 | Boyd, Steven | ** Realignment work schedule |
34 | 13 | Boyd, Steven | ** Tracker patch panel leak work |
35 | 10 | Boyd, Steven | ** Date of DSA closure and work required before hand |
36 | 3 | Boyd, Steven | ** "Henry's January workplan":http://micewww.pp.rl.ac.uk/projects/operations/wiki/WorkPlanJanuary2016 |
37 | 1 | Boyd, Steven | ** "Daresbury workplan":http://micewww.pp.rl.ac.uk/attachments/5602/MICE_Work_List_18-01-16_v2_1.pdf |
38 | * AOB |
||
39 | * Date of next meeting : February 2nd |
||
40 | |||
41 | |||
42 | h2. Minutes |
||
43 | 15 | Boyd, Steven | |
44 | Present: Andy, Vicky Bayliss, Paolo, Durga, Chris R, Steve B, Paul K, Trevor H, Jaroslaw, Mike C, Victoria, John N, Paul H, Henry N, Ken L |
||
45 | |||
46 | h3. Review of action items |
||
47 | |||
48 | * Realignment work |
||
49 | ** *Action Item (Colin to prod John C)* : need to get John's data with fiducials in it to get to Tony Middleton. _Expected yesterday_ |
||
50 | 18 | Boyd, Steven | *** _Done_ |
51 | 15 | Boyd, Steven | * Simulation of material in the interspace |
52 | ** *Action Item* : John to do a simulation with air in the interspace and He in the absorber volume |
||
53 | 18 | Boyd, Steven | *** _John presented plots to show that there is little effect. However, these plots include systematic effects in the reconstruction |
54 | that are being worked on._ |
||
55 | *** _*Action Item* : John will look at the effect of air on the MC truth scattering angle_ |
||
56 | 15 | Boyd, Steven | ** *Action Item* : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded |
57 | 18 | Boyd, Steven | *** _*Action Item* : DC to work with Pierrick to ensure this_ |
58 | 15 | Boyd, Steven | ** *Action Item* : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum |
59 | 18 | Boyd, Steven | *** _Alan says that the LiH absorber can withstand vacuum._ |
60 | 15 | Boyd, Steven | * Cooling water discussion |
61 | ** *Action Item* : Colin and John to talk and agree on the plan. |
||
62 | 18 | Boyd, Steven | *** _Discussion complete but further work needed_ |
63 | 15 | Boyd, Steven | * ISIS Monitoring of SS magnets |
64 | ** *Action Item* : Josef/Colin to check with Alan about monitoring requirements for the SS magnets. |
||
65 | 18 | Boyd, Steven | *** _Josef doesn't think we need to monitor to the level we are doing now. Proposed to stop the 4-hour monitoring and just leave |
66 | the buzzer in the control room._ |
||
67 | 17 | Boyd, Steven | *** *Action Item* : Steve to talk to Alan Stevens about this possibility- |
68 | 15 | Boyd, Steven | * BLOCS and SOCS |
69 | ** *Action Item* : Steve to remind Adam that we need SOCs. |
||
70 | 18 | Boyd, Steven | *** _Done_ |
71 | 15 | Boyd, Steven | * DS PS Installation |
72 | ** *Action Item* : Colin is to get an estimate for power supply installation time. |
||
73 | 18 | Boyd, Steven | *** _2 sets of 2-3 days over a space of 2 weeks._ |
74 | 15 | Boyd, Steven | |
75 | h3. Updates and future schedule |
||
76 | |||
77 | 20 | Boyd, Steven | *DS Power Supply work |
78 | ** (Trevor) The PS is expected to return in mid-February (8th - 15th). He plans to start commissioning |
||
79 | the new PS on Feb 22nd. This could roll into the next week (which is nominally a data-taking week so may need to be postponed). |
||
80 | ** (Vicky) The compressor has a leak and will be fixed on Friday 29th and cleaned on Monday 1st |
||
81 | ** (Discussion between Mike, Vicky and Henry) : The Linde IOC has been beahving oddly and we need to understand it. |
||
82 | *** *Action Item* : Paolo, Mike and Vicky will investigate this issue ASAP. We can't cool down magnet until this is understood. |
||
83 | ** (Discussion between Mike, Vicky and Henry) : We plan to start the cooldown of the cold box on Tuesday 2nd, with work on the controls and |
||
84 | monitoring doable whilst that is happening. If all is well, we can start the cooldown of the magnet on Thursday 4th. |
||
85 | ** (Discussion between Paolo and Mike) : we need a way of monitoring the status of the Linde IOC machine |
||
86 | ** *Action Item* : Paolo/Pierrick to consider this issue |
||
87 | |||
88 | * Controls and Monitoring : Network reconfiguration |
||
89 | ** (Paolo) Most work has been done but we wait for Pierrick to return to finish it. There is potentially another week of work with 1 - 2 |
||
90 | days of severe disruption to the network whilst things are tested. This needs to be advertised well in advance |
||
91 | 1 | Boyd, Steven | |
92 | 20 | Boyd, Steven | * Realignment work |
93 | ** Drilling of holes in the flanges will be done by Thursday 28th |
||
94 | ** Template to show where the holes should go have been designed but need to be constructed. |
||
95 | ** EMR will not be moved further at this time |
||
96 | ** Revisit next week |
||
97 | |||
98 | * Tracker patch panel leak work |
||
99 | ** John wants to take the front of the PRY off to get at the tracker leak, which he may have a solution for. This can be done now. |
||
100 | |||
101 | * DSA Work before closure on Friday 29th |
||
102 | ** Need to reconnect the dipole magnets, and test all magnets in the DSA |
||
103 | ** New drip trays will not be installed at this point |
||
104 | |||
105 | * Electrical work schedule |
||
106 | ** (Trevor) Trevor/Steve Griffiths will be here on Monday 1st/Tuesday 2nd to do a low current test on the SS. This will |
||
107 | require the Hall closed up and the SS permit to be granted. DC will work with Trevor on this. |
||
108 | *** *Action Item* : (John G) need to put barriers around any open part of the SS |
||
109 | *** *Action Item* : Andy/DC will organise the magnetic sweep |
||
110 | ** A long discussion on the Daresbury scheduled ensued. We decided we needed a longer discussion. |
||
111 | ** *Action Item* : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time |
||
112 | and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th. |